OpenStudy (anonymous):

Points D(-4, 6), E(5, 3), and F(3, -2) are the vertices of triangle DEF. Find the perimeter of the triangle.

OpenStudy (anonymous):

you need to use pythagorus' theorem, a^2=b^2=c^2 you need to create right angled triangles around the triangle DEF to find all of the sides, DF, EF and DE can be made to be hypotoneuses of three seperate right angled triangles, you will end up with an equation like.. \[\sqrt{DF}+\sqrt{EF}+\sqrt{DE}= P\]